OpenStudy (anonymous):

5(2x - 7) = 33 - 3(6x + 4)

OpenStudy (anonymous):

comes to 10x-35=33-18x-12 which is 28x=56 so x=56/28 or x=2

OpenStudy (anonymous):

Can you show me in steps please

OpenStudy (gw2011):

5(2x-7)=33-3(6x+4) 10x-35=33-18x-12 10x+18x=33-12+35 28x=56 x=2